How do I solve the following by substitution method? y=-4x+8 , y=x+7
Answer by checkley75(3666)

y=x+7 now substitute (x+7) for y in the other equatuion and solve for x
x+7=-4x+8
x+4x=8-7
5x=1
x=1/5 answer.
now substitute 1/5 for x & solve for y
y=1/5+7
y=7 1/5 or 36/5 answer.
proof
36/5=-4*1/5+8
36/5=-4/5+8
35/5=(-4+40)/5
36/5=36/5